plz explain this question in detail

Q8. Two wave sources, separated by a distance of 9.4 m, emit in phase sound of wavelength
6.8 m. Destructive interference will cause the formation of nodes on the line between the
sources at the following distances from one of the sources:
A. 1.85 m only
B. 3.133 m only
C. 1.3 m, 4.7 m and 8.1 m
D. 3.0 m and 6.4 m
E. 4.7 m only
Answer #1

for destructive interference,

x = \lambda(n+1/2) (where x is the path difference, \lambda = 6.8, n = 0,1,2,3.....)

on a line between the two sources let the point of destructive interference be at a distance of a from one source,

then the path difference = 9.4-a-a = 9.4-2a

-->  9.4-2a = \lambda(n+1/2)

for n=0

a=6.4m

for n=1

a=3m

thus the correct option is D
Answer #2

the answer is E

4.7 metres only.

Think of it in this manner...

Both waves would come from opposite directions.

Since they are in phase, their speeds would be equal.

They would meet each other at the centre.

At the centre, when they meet, due to opposite directions, they will cancel each other.

thus, a node would be formed. ASK IF ANY DOUBTS...

THANK YOU. DO RATE
Answer #3

B.3.133 m only